I need help with this word problem, I've tried everything, but can't get a real answer:
The altitude of a triangle is 8cm longer than the base. If the area of a triangle is 24cm, find the altitude.

Answer by checkley71(8403)

area=bh/2
h=b+8
area=b(b+8)/2
24=(b^2+8b)/2 now cross multiply.
b^2+8b=48
b^2+8b-48=0
(b+12)(b-4)=0
b-4=0
b=4 answer for the base.
h=4+8=12 answer for the height.
proof
24=4*12/2
24=48/2
24=24